OpenStudy (anonymous):

related rates problem need help stat! please!! Two cars start moving from the same poin. one travels south at 60mi/h and the other travels west at 25 mi/h. at what rate is the distance between the cars increasing two hours later?

OpenStudy (anonymous):

Any help on this problem would be greatly appreciated!

OpenStudy (amistre64):

looks like a right triangle relationship to me ... how would we write this up?

OpenStudy (amistre64):

what formula do we know that compares and relates the sides of a right triangle?

OpenStudy (anonymous):

Well I think we would use implicit differentiation but would we use a formula for right triangles?

OpenStudy (anonymous):

Do we use pythagorean theorem a^2 +b^2= c^2?

OpenStudy (amistre64):



OpenStudy (amistre64):

pythag is good what is the derivative of: d^2 = w^2 + s^2 ?

OpenStudy (anonymous):

i was going to say that 2 is a constant so it equals 0. but that seems incorrect would it be 2d because of the power rule? I'm not quite sure.

OpenStudy (amistre64):

d^2 derives to 2d d' and the others follow suit giving us: 2d d' = 2w w' + 2s s' we can cancel out the 2s and solve for d' d' = (w w' + s s')/d all these parts are known or can be easily determined now s'=60, w'=25, at 2 hours, w=60(2) and s=25(2), and we can know d with pythag

OpenStudy (amistre64):

lol, um, cancel out all the 2 parts, not "2s" :)

OpenStudy (anonymous):

oh okay thank you once the two parts are cancelled out are you left with w=120 and s= 50 then take the square root or no?

OpenStudy (anonymous):

like what is the format of the problem from this particular point?

OpenStudy (amistre64):

d = sqrt(120^2 + 50^2) yes then you have all the parts for the d' rate of change

OpenStudy (amistre64):

\[d'=\frac{w'w+s's}{d}\] \[d'=\frac{60(120)+25(50)}{\sqrt{120^2+50^2}}\]

OpenStudy (anonymous):

ThanK you!!!

OpenStudy (amistre64):

youre welcome